# Jin Li    Ye Luo    Xiaowei Zhang

### v3

@@ Appendix B Proofs for Multivariate Normal Distribution
Appendix C Proofs for Bivariate $t$ Distribution
@@
$\mathbb{E}[U|U>z]=\mu+\sigma\frac{\delta\Gamma\left(\frac{\delta+1}{2}\right)}{(\delta-1)\sqrt{\delta\pi}\Gamma\left(\frac{\delta}{2}\right)}\left(1+\frac{(z-\mu)^{2}}{\sigma^{2}\delta}\right)^{-\frac{\delta-1}{2}}\frac{1}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)},$ (C.1)
@@
where the last step follows from the gamma function’s property $\Gamma(z+1)=z\Gamma(z)$. Hence, we can rewrite (C.1) as
$\mathbb{E}[U|U>z]=\mu+\sigma\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)}\right).$ (C.2)
@@
where the third equality holds because $\mathbb{E}[U+V|U>z]=\mathbb{E}[U+V|V>z]$ (symmetry), and the last equality follows from (C.2) and (C.3). ∎
@@
Let $X$ be a standard univariate $t$ distribution. Then, (C.2) implies that
@@
for all $\alpha>0$. Setting $(z^{*}-\mu)/\sigma=\alpha$, which is positive by the definition of $z^{*}$, in (C.4) and using the assumption $\mu<0$, we conclude that $f(z^{*})>0$. ∎

### v4

@@ B.2 Proofs for Multivariate Normal Distribution
B.3 Proofs for Bivariate $t$ Distribution
@@
$\mathbb{E}[U|U>z]=\mu+\sigma\frac{\delta\Gamma\left(\frac{\delta+1}{2}\right)}{(\delta-1)\sqrt{\delta\pi}\Gamma\left(\frac{\delta}{2}\right)}\left(1+\frac{(z-\mu)^{2}}{\sigma^{2}\delta}\right)^{-\frac{\delta-1}{2}}\frac{1}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)},$ (B.3)
@@
where the last step follows from the gamma function’s property $\Gamma(z+1)=z\Gamma(z)$. Hence, we can rewrite (B.3) as
$\mathbb{E}[U|U>z]=\mu+\sigma\left(\frac{\delta}{\delta-2}\right)\left(\frac{t_{\delta-2}\left(\frac{z-\mu}{\sigma};0,\frac{\delta}{\delta-2}\right)}{1-T_{\delta}\left(\frac{z-\mu}{\sigma};0,1\right)}\right).$ (B.4)
@@
where the third equality holds because $\mathbb{E}[U+V|U>z]=\mathbb{E}[U+V|V>z]$ (symmetry), and the last equality follows from (B.4) and (B.5). ∎
@@
Let $X$ be a standard univariate $t$ distribution. Then, (B.4) implies that
@@
for all $\alpha>0$. Setting $(z^{*}-\mu)/\sigma=\alpha$, which is positive by the definition of $z^{*}$, in (B.6) and using the assumption $\mu<0$, we conclude that $f(z^{*})>0$. ∎